Please explain how to determine polarity for me.

Rank the following covalent bonds by decreasing polarity. (Most polar = 1, Least polar = 4)

S-Cl
S-H
S-O
S-F
Answer #1

Ans :

The polarity of the bonds depends on the electronegativity difference between the constituent atoms , as in larger the difference , higher the polarity.

Here S is common in all the bonds and the other atom changes :

S-Cl = difference : 0.5

S-H = difference : 0.4

S-O = difference : 1

S - F = difference : 1.5

Now here , ranking :

S-Cl : 3

S-H : 4 ( least polar)

S-O : 2

S-F : 1 ( most polar)